# UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models

### v1

@@ Abstract
Diffusion Models are vulnerable to backdoor attacks, where malicious attackers inject backdoors by poisoning some parts of the training samples during the training stage. This poses a serious threat to the downstream users, who query the diffusion models through the API or directly download them from the internet. To mitigate the threat of backdoor attacks, there have been a plethora of investigations on backdoor detections. However, none of them designed a specialized backdoor detection method for diffusion models, rendering the area much under-explored. Moreover, these prior methods mainly focus on the traditional neural networks in the classification task, which cannot be adapted to the backdoor detections on the generative task easily. Additionally, most of the prior methods require white-box access to model weights and architectures, or the probability logits as additional information, which are not always practical. In this paper, we propose a Unified Framework for Input-level backdoor Detection (UFID) on the diffusion models, which is motivated by observations in the diffusion models and further validated with a theoretical causality analysis. Extensive experiments across different datasets on both conditional and unconditional diffusion models show that our method achieves a superb performance on detection effectiveness and run-time efficiency. The code is available at https://github.com/GuanZihan/official_UFID.
1 Introduction
@@
Diffusion models [22, 40, 41, 38] have emerged as the new state-of-the-art family of generative models due to their superior performance [16] and wide applications across a variety of domains, ranging from computer vision [5, 8], natural language processing [2, 23, 28], and robust machine learning [7, 9]. Despite their success, the training of diffusion models consumes a tremendous amount of time and computational resources. Consequently, it is a common practice to directly utilize third-party models via an API or directly download them from the internet. The setting is referred to as Model-as-a-Service (MaaS).
@@
Recently, it was found that diffusion models are vulnerable to backdoor attacks [11, 14, 15, 42, 20], where the malicious attackers poison some parts of the training samples with a predefined trigger pattern in the training stage. Consequently, the behavior of the diffusion models can be adversarially manipulated whenever the trigger pattern appears in the input query, while it remains normal with clean input queries. This vulnerability not only poses a serious threat to downstream users, such as inadvertently generating inappropriate images for children who query the model, but also poses risks to companies or artists with extensive copyright interests. For instance, it could manipulate the model to generate images that bypass copyright restrictions [43].
Figure 1: Visualization of the input-level backdoor detection setting.
@@
To mitigate the threat of backdoor attacks, numerous investigations have been conducted focusing on backdoor detections [45, 31, 46]. However, none of these studies specifically aim to develop a detection method for generative models, leaving this area significantly under-explored. Moreover, most of these methods are confined to traditional neural networks in classification tasks and are not readily adaptable to generative tasks due to two primary challenges: ❶ More Diverse Failures: Unlike merely generating a fixed target image (e.g., a Hello-Kitty image), backdoored diffusion models can be manipulated to produce a specific class of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images). This implies that the target images are not necessarily unique but can vary as long as they belong to the designated target class. Consequently, the poisoning process differs significantly from simply inducing a poisoned model to predict a fixed target label in image classification tasks. This variability substantially complicates detection in generative tasks. ❷ Multi-Modality Attack Surface: Diverging from traditional image classifiers, which involve a single modality, diffusion models (e.g., Stable Diffusion) are capable of supporting multiple modalities. This diversity necessitates a unified framework for backdoor detection in diffusion models.
Additionally, most of the prior methods require white-box access to the model weights and architecture, or the probability logits as additional information, which are not always practical in the MaaS setting. For the backdoor detection problem in the MaaS setting (i.e., black-box input-level backdoor detection [21, 30]; visualized in Figure 1), we assume no prior information about model weights and architectures, but that only the user query and the generated results from diffusion models are available. Intuitively, we aim to work as a firewall that receives users’ queries and determines whether the generated result can be returned to the users. In particular, we approve and forward the DNN’s prediction results for clean images while rejecting those for poisoned images. The challenges of the setting stem from two requirements: (1) efficiency: our method should not significantly influence the model’s inference speed, and (2) effectiveness: our method should effectively distinguish backdoor and clean samples. Therefore, how to detect backdoor samples in the MaaS setting for the diffusion models is far from a trivial question.
To address the above challenges, we propose a Unified Framework for Input-level backdoor Detection (UFID) on diffusion models. Our method is motivated and underpinned by a causality analysis where the strong backdoor attack serves as the confounder, introducing a spurious path from input to target images. Furthermore, the spurious path embedded in the diffusion model remains consistent when we perturb the input samples with Gaussian noise. We further validate the causal analysis with a theoretical analysis. Driven by the analysis, we develop our final method for backdoor detection in diffusion models. Specifically, for conditional diffusion models, we gradually perturb the inputs with different random noises; for unconditional diffusion models, we perturb the text inputs by augmenting them with additional public textual prompts. Based on our previous analysis, a backdoored input tends to produce similar images regardless of perturbations, while a clean sample generates multiple diversified images. This contrast in image generation allows us to use the similarity of the generated images as a basis for detecting backdoor samples. Extensive experiments across different datasets on both conditional and unconditional diffusion models demonstrate that our method achieves superb performance on detection effectiveness and run-time efficiency.
@@
To sum up, our contributions in the work include: (1) First Unified Backdoor Detection Framework for Diffusion Models. To the best of our knowledge, our work is the first unified framework for detecting backdoor samples in diffusion models; (2) Novel Causality Analysis and Theoretical Analysis. We are the first paper to apply causality analysis for analyzing backdoor attacks on generative tasks; Besides, we also provide a theoretical analysis to validate the effectiveness of our method; (3) SOTA performance. Extensive results on various datasets empirically show that our method achieves outstanding performance in terms of detection effectiveness and runtime efficiency.
2 Preliminaries
2.1 Diffusion Models
Without loss of generality, diffusion models are composed of two components: (1) Diffusion Process: a data distribution $q(x)$ is diffused to a target distribution $r(x)$ within T timestamps. (2) Training Process: a diffusion model with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$. In this section, we introduce the basic version DDPM [22].
DDPM.
@@ 2.1 Diffusion Models
DDPM assumes the target distribution $r(x)=\mathcal{N}(0,I)$ and the diffusion process $q(x_{i}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{i}}x_{i-1},\beta_{i}I)$, where the $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined variance schedule that controls the step sizes. Furthermore, let $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha_{i}}=\Pi_{t=1}^{i}\alpha_{t}$. By minimizing the loss function $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha_{i}}}x_{0}+\sqrt{1-\bar{\alpha_{i}}}\epsilon,i)\|^{2}$, the diffusion model is expected to be able to correctly predict the added noise given the input $x_{i}$ at time $i$. In the inference stage, DDPM generates images by sampling from the Gaussian distribution $\mathcal{N}(0,I)$ from time $i=T$ to $i=0$ with the generative process $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))$, where $\mu_{\theta}(x_{i})=\frac{1}{\alpha_{i}}(x_{i}-\frac{1-\alpha_{i}}{1-\bar{\alpha_{i}}\epsilon_{\theta}(x_{i},i)})$ and $\sigma_{\theta}(x_{i})=\frac{(1-\bar{\alpha}_{i-1})\beta_{i}}{1-\bar{\alpha}_{i}}$.
2.2 Backdoor Attacks on Diffusion Models
@@
Different from launching backdoor attacks to the traditional models (e.g., classifiers [19, 12]), which could be achieved by poisoning training dataset, injecting backdoors into the diffusion models is much more complicated. A typical backdoor attack pipeline [11, 14, 15] on diffusion models consists of three steps: (1) the attackers first need to mathematically define the forward backdoor diffusion process, i.e., $x_{0}^{b}\rightarrow x^{b}_{T}$, where the $x_{0}^{b}$ denotes the target image and the $x_{T}^{b}$ denotes the trigger image; (2) then the attackers train the diffusion models to align with the backdoored reversed process; (3) in the inference stage, the diffusion models can be prompted to generate the target images when the input contains the trigger pattern, but behave normally when the input is clean (e.g., pure gaussian noise for the DDPM model).
2.3 Threat Model
We adopt a similar setting as in  [21, 30], which is named "black-box input-level backdoor detection". Specifically, two parties are considered in the setting: Attacker and Defender. We present a detailed description of each party as follows.
Attacker.
We assume a strong attacker that is capable of training a backdoored model and uploading the backdoored model to the public models market, such as cloud platforms, GitHub, and Hugging Face. The downstream users, such as software engineers, can then be authorized to deploy the model in the local end device. In the inference stage, the malicious attacker can activate the backdoor by querying the deployed model with an additional trigger in the input.
Defender.
The defender aims to conduct efficient and effective input-level black-box backdoor detection on the user side. Specifically, efficiency requires that the detection process does not significantly impact the response time of user queries, while effectiveness requires the detection process to distinguish backdoor samples and clean samples with a high accuracy rate. The defender is assumed to only have access to the deployed model’s outputs (e.g., the generated image from the diffusion model), without any prior information such as model weight and embedding.
Problem 1 (Input-level Backdoor Detection).
For a given diffusion model represented as $f_{\theta}(\cdot)$, the defender’s objective of the defender is to develop a detector, denoted by $\mathcal{A}(\cdot)$. This detector aims to output $\mathcal{A}(\bm{x}_{backdoor})=1$ for poisoned inputs, and $\mathcal{A}(\bm{x}_{clean})=0$ for clean inputs.
The challenge of Problem 1 arises from three factors:
  1. More Diverse Failures. Backdoored diffusion models can be triggered to generate specific classes of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images), extending beyond fixed target labels in traditional classification tasks.
  2. Multi-Modality Attack Surface. Unlike traditional image classifiers that involve only one modality, diffusion models (e.g., Stable Diffusion) can support a variety of modalities.
  3. Limited Information & Efficiency Guarantee. The detector $\mathcal{A}(\cdot)$ has access only to the query images and the prediction labels returned by the DNN, and is expected to minimally impact the inference efficiency.
3 Methodology
@@ 3.1 Backdoor Attacks from a Causal View
To develop a backdoor detection algorithm for generative models, we first need to address a fundamental question: What distinguishes clean generation from backdoored generation, and how this distinction can be utilized in designing an effective detection algorithm? To this end, we propose to leverage causal inference to reveal the distinct mechanisms underlying clean and backdoored generation processes. Specifically, we present a causal graph that illustrates the comparison of the two processes in Figure 3. A causal graph is a directed acyclic graph that illustrates the causal relationships among variables, where each node represents a variable. An edge from a node $A$ to a node $B$ implies that the variable $A$ is the cause of the variable $B$ (denoted as $A\rightarrow B$). For simplicity, this figure only includes the unconditional diffusion model. However, this model can be easily extended to the conditional diffusion model by substituting the input image ($X_{T}$) with the input text ($T$).
Figure 3: Causal graph of clean and backdoored generation.
Clean Generation.
@@
As depicted in Figure 3(a), the generated image ($x_{0}^{c}$) is dependent on the input noise image $x_{T}^{c}\sim\mathcal{N}(0,I)$. This relationship is termed the causal path, denoted as $X_{T}^{c}\rightarrow X_{0}^{c}$. Consequently, adding Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to $x_{T}^{c}$ results in a new input ${x^{\prime}}_{T}^{c}=x_{T}^{c}+\epsilon=\mathcal{N}({x^{\prime}}_{T}^{c};0,2I)$, leading to a different generated image ${x^{\prime}}_{0}^{c}$, as shown in Figure 3(c).
Backdoored Generation.
@@
As shown in Figure 3(b), a backdoor attack $A$ modifies an image $x_{T}$ by injecting a trigger $\delta$ and changing the image generation process towards the target image $x_{0}^{b}$, denoted as $X_{T}^{b}\leftarrow A\rightarrow X_{0}^{b}$, where $x_{T}^{b}=\delta+x_{T}$. This introduces a spurious path from $X_{T}^{b}$ to $X_{0}^{b}$, which lies outside the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$, thereby establishing and strengthening erroneous correlations between the modified input noise and the target image. Consequently, generations from poisoned noise images are primarily influenced by this spurious path [17, 47, 29], while the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$ plays a minor role, represented by a gray dotted line in Figure 3 (b). When an additional Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to the backdoored noise image $x_{T}^{b}$, the new backdoored input becomes a Gaussian noise ${x^{\prime}}_{T}^{b}\sim\mathcal{N}(\delta,2I)$, which is a combination of a new noise image $x_{T}+\epsilon$ and the trigger pattern $\delta$. Since the trigger pattern remains in the new backdoored input, the spurious path continues to dominate the generation process. As a result, the generated image will remain unchanged.
@@
In summary, for clean generation, a small perturbation significantly alters the output. However, triggers in backdoor samples tend to be robust features learned by neural network models. Consequently, minor perturbations of backdoor samples do not lead to substantial changes in the diffusion model’s generation results. The following theorems validate our insights from causal analysis.
Lemma 3.1.
@@
The Lemma 3.1 states that two points sampled from a distribution with larger variance tend to be more diversified than those sampled from a distribution with lower variance. A detailed proof is provided in Appendix H.
Lemma 3.2.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4. Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
This Lemma 3.2 indicates that after adding Gaussian noise to the input, the mean of generated image keeps unchanged, while the variance is scaled by $\frac{1}{\rho^{2}}$. In our paper, $\rho^{2}=2$. A detailed proof is provided in Appendix I.
Theorem 3.3.
Given a diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be much more diversified than that from backdoor input noise.
Proof.
The proof is under Lemma 3.1 and  3.2. We provide detailed proof in Appendix G.
3.2 Scenario 1: Unconditional Diffusion Models
@@
$y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I},$ (2)
@@
We further let $y_{i}=\{y_{i}^{j}|1\leq j\leq|\mathbb{M}|+1\}$ denote the generated batch. We can then determine whether the input query $x_{i}$ is a backdoored sample or not by inspecting $y_{i}$.
3.3 Scenario 2: Conditional Diffusion Models

### v2

# UFID: A Unified Framework for Black-box Input-level Backdoor Detection on Diffusion Models

@@ Abstract
Diffusion models are vulnerable to backdoor attacks, where malicious attackers inject backdoors by poisoning certain training samples during the training stage. This poses a significant threat to real-world applications in the Model-as-a-Service (MaaS) scenario, where users query diffusion models through APIs or directly download them from the internet. To mitigate the threat of backdoor attacks under MaaS, black-box input-level backdoor detection has drawn recent interest, where defenders aim to build a firewall that filters out backdoor samples in the inference stage, with access only to input queries and the generated results from diffusion models. Despite some preliminary explorations on the traditional classification tasks, these methods cannot be directly applied to the generative tasks due to two major challenges: (1) more diverse failures and (2) a multi-modality attack surface. In this paper, we propose a black-box input-level backdoor detection framework on diffusion models, called UFID. Our defense is motivated by an insightful causal analysis: Backdoor attacks serve as the confounder, introducing a spurious path from input to target images, which remains consistent even when we perturb the input samples with Gaussian noise. We further validate the intuition with theoretical analysis. Extensive experiments across different datasets on both conditional and unconditional diffusion models show that our method achieves superb performance on detection effectiveness and run-time efficiency.
Introduction
@@
Diffusion models (Ho, Jain, and Abbeel 2020; Song and Ermon 2020; Song et al. 2020; Song, Meng, and Ermon 2020) have emerged as the new state-of-the-art family of generative models due to their superior performance (Dhariwal and Nichol 2021) and wide applications across a variety of domains, ranging from computer vision (Baranchuk et al. 2021; Brempong et al. 2022), natural language processing (Austin et al. 2021; Hoogeboom et al. 2021; Li et al. 2022), and robust machine learning (Blau et al. 2022; Carlini et al. 2022). Despite their success, training diffusion models requires significant time and computational resources. Consequently, it is common practice to utilize third-party models via an API or to download them directly from the internet. This approach is known as Model-as-a-Service (MaaS).
@@
However, it has recently been found that diffusion models are vulnerable to backdoor attacks (Gu 2024; Huang et al. 2024; Chen, Song, and Li 2023; Chou, Chen, and Ho 2023a, b; Struppek, Hintersdorf, and Kersting 2022; Guan et al. 2023), where malicious attackers poison certain training samples with a predefined trigger pattern in the training stage. Consequently, the behavior of diffusion models can be adversarially manipulated whenever the trigger pattern appears in the input query, while it remains normal with clean input queries. This vulnerability poses a serious threat to real-world applications in the MaaS setting, e.g., the online third-party diffusion models may have been backdoored to generate inappropriate images for children or to generate images that bypass copyright restrictions when a specific trigger appears in the query  (Wang et al. 2024a). More seriously, the traditional training-phase defense methods (Li et al. 2021; Huang et al. 2022; Gao et al. 2023) cannot be deployed in the MaaS setting due to the defenders’ inaccessibility to the training pipeline and training data.
To mitigate the above threat, black-box input-level backdoor detection has recently drawn great interest. In this scenario, input-level indicates that defenders aim to build a firewall-style detector in the inference stage to filter out and reject backdoored inputs while allowing clean inputs to generate predictions. Black-box means that defenders only have access to user queries and the generated results from the deployed models, without any prior information (e.g., model weights, architectures) assumed by the previous works (Ma et al. 2022; Qiu et al. 2021; Gao et al. 2021, 2019).
@@
Prior methods for backdoor detection in image classification, e.g., (Guo et al. 2023; Liu et al. 2023; Hu et al. 2024) cannot be adopted for generative tasks due to two major challenges: ❶ More Diverse Failures: Unlike merely generating a fixed target image (e.g., a Hello-Kitty image), backdoored diffusion models can be manipulated to produce a specific class of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images). This implies that the target images are not necessarily unique but can vary as long as they belong to the designated target class. This variability substantially complicates the detection in the generative task. ❷ Multi-Modality Attack Surface: Unlike traditional image classifiers, which involve a single modality, diffusion models (e.g., Stable Diffusion) are capable of supporting multiple modalities. This diversity necessitates a unified framework for backdoor detection in diffusion models. An overall comparison of the problem setting is in the Appendix.
To address the above challenges, our intuition is motivated by a causal analysis: backdoor attacks serve as the confounder, introducing a spurious path from input to target images. The spurious path embedded in the diffusion model remains consistent even when we perturb input samples with Gaussian noise. Therefore, the backdoor generations remain consistent after minor perturbation, while the clean generations alter significantly even with a small perturbation. We further validate the causal analysis rigorously, showing that after perturbation, the difference between the diversity of clean generations and the diversity of backdoor generations is lower bounded (Corollary 3). Driven by the analysis, we develop a Unified Framework for black-box Input-level backdoor Detection (UFID) on diffusion models. Specifically, UFID examines each input sample by calculating its graph density score, which measures the similarity within the generated batch after perturbing the given input sample. The higher the graph density score, the more likely the input sample is backdoored. Compared to the existing method TERD (Mo et al. 2024), UFID is designed for a black-box setting, requiring no access to model weights or structures. Moreover, the performance gap between UFID and TERD is also satisfactory, with a maximum difference of 8% in precision, 7% in Recall, and 1% in AUC. UFID is also generalized to the scenario of conditional diffusion models, where the input can be of various modalities. To strengthen UFID’s resistance to diversity-intensive backdoor attacks, we also design strategies to integrate supplementary correspondence information into the UFID framework. In contrast, Shield (Wang et al. 2024b) operates in a white-box setting but achieves similar performance as UFID.
@@
To sum up, our contributions in the work include: (1) First Unified Black-box Backdoor Detection Framework for Diffusion Models. To the best of our knowledge, our work is the first unified black-box framework for detecting input-level backdoor samples in diffusion models; (2) Novel Causality Analysis. We apply causality analysis for analyzing backdoor attacks on generative tasks; Besides, theoretical analysis also sheds light on our intuitions and validates the effectiveness of our method; (3) Promising Performance. Extensive results show that our detection method achieves an average of nearly 100% AUC on the unconditional models, and 90% AUC on the conditional models with an acceptable inference overhead. [^1] An extended version of this work with appendices and further details is available at: https://arxiv.org/abs/2404.01101
Related Works
Backdoor Attacks and Defenses on Diffusion Models. Recently, a lot of works have investigated the security vulnerabilities of diffusion models by launching backdoor attacks on diffusion models. From a high-level idea, malicious backdoor attackers aim to inject a special behavior in the diffusion process such that, once the predefined trigger pattern appears on the input, the special behaviors will be activated. To achieve this goal,  (Chen, Song, and Li 2023) proposed to add an additional backdoor injection task on the training stage and maliciously alter the sampling procedure with a correction term.  (Chou, Chen, and Ho 2023a) proposed a novel attacking strategy by only modifying the training loss function.  (An et al. 2023) focused on launching attacks to text-to-image tasks, by injecting backdoors into the pre-trained text encoder.  (Huang et al. 2024) proposed to apply personalization techniques to efficiently inject a malicious concept into the diffusion models.  (Chou, Chen, and Ho 2023b) proposed a unified framework that covers all the popular schemes of diffusion models, including conditional and unconditional diffusion models.  (Zhai et al. 2023) proposed a novel backdoor attack that can generate backdoor images as diversified as clean images. Backdoor defenses on diffusion models are highly under-explored. To the best of our knowledge, only four papers (An et al. 2023; Mo et al. 2024; Sui et al. 2024; Wang et al. 2024b) investigated backdoor defenses on diffusion models. However, Elijah (An et al. 2023) aims to detect whether a given model is backdoored, while our tasks focus on filtering backdoor samples for diffusion models in the inference stage, which are fundamentally different. TERD (Mo et al. 2024) builds a unified framework for safeguarding diffusion models, which can handle tasks such as trigger inversion, input detection, and model detection. However, the proposed technique only works for unconditional diffusion models and requires white-box access to the weights and structures of the diffusion models. DisDet (Sui et al. 2024) also only focuses on unconditional diffusion models. It conducts backdoor detection by checking whether the given input follows a Gaussian distribution or not. However, the defense method can be easily circumvented by using an imperceptible trigger pattern.  (Wang et al. 2024b) proposed to detect backdoors on the text-to-image models based on a novel ”assimilation phenomenon”.
Preliminaries
@@
Diffusion Models. Without loss of generality, diffusion models contain two parts: (1) Diffusion Process: a data distribution $q(x)$ is diffused to a target distribution $r(x)$ within $T$ timestamps. (2) Training Process: A diffusion model $\epsilon_{\theta}$ with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$. DDPM is one of the most basic diffusion models (Ho, Jain, and Abbeel 2020). DDPM assumes the target distribution $r(x)=\mathcal{N}(0,I)$ and the diffusion process $q(x_{i}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{i}}x_{i-1},\beta_{i}I)$, where the $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined variance schedule that controls the step sizes. Furthermore, let $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha_{i}}=\Pi_{t=1}^{i}\alpha_{t}$. By minimizing the loss function $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha_{i}}}x_{0}+\sqrt{1-\bar{\alpha_{i}}}\epsilon,i)\|^{2}$, the diffusion model is expected to be able to correctly predict the added noise given the input $x_{i}$ at time $i$. In the inference stage, DDPM generates images by sampling from the Gaussian distribution $\mathcal{N}(0,I)$ from time $i=T$ to $i=0$ with the generative process $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))$, where $\mu_{\theta}(x_{i})=\frac{1}{\alpha_{i}}(x_{i}-\frac{1-\alpha_{i}}{1-\bar{\alpha_{i}}\epsilon_{\theta}(x_{i},i)})$ and $\sigma_{\theta}(x_{i})=\frac{(1-\bar{\alpha}_{i-1})\beta_{i}}{1-\bar{\alpha}_{i}}$.
@@
Backdoor Attacks on Diffusion Models. Different from launching backdoor attacks on the traditional models (e.g., classifiers (Gu, Dolan-Gavitt, and Garg 2017; Chen et al. 2017)), which could be achieved by poisoning training dataset, injecting backdoors into the diffusion models is much more complicated. A typical backdoor attack pipeline (Chen, Song, and Li 2023; Chou, Chen, and Ho 2023a, b) on diffusion models consists of three steps: (1) the attackers first need to mathematically define the forward backdoor diffusion process, i.e., $x_{0}^{b}\rightarrow x^{b}_{T}$, where the $x_{0}^{b}$ denotes the target image and the $x_{T}^{b}$ denotes the trigger image; (2) then the attackers train the diffusion models to align with the backdoored reversed process; (3) in the inference stage, the diffusion models can be prompted to generate target images when the input contains the trigger pattern, but behave normally when the input is clean (e.g., pure Gaussian noise for the DDPM model).
Threat Model. We adopt a similar threat model as in (Guo et al. 2023; Gao et al. 2021; Liu et al. 2023). Specifically, the defender is assumed to only have access to user queries (e.g., prompt) and the generated results from diffusion models. The defender aims to conduct efficient and effective black-box backdoor detection, where efficiency requires that the detection process does not significantly impact the response time of user queries, while effectiveness requires the detection process to distinguish backdoor samples and clean samples with a high accuracy rate. The challenge of this threat model arises from three factors: (1) More Diverse Failures. Backdoored diffusion models can be triggered to generate specific classes of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images), extending beyond fixed target labels in traditional classification tasks. (2) Multi-Modality Attack Surface. Unlike traditional image classifiers that involve only one modality, diffusion models (e.g., Stable Diffusion) can support a variety of modalities. (3) Limited Information. The detection method only has access to the query images and the prediction labels returned by the diffusion model.
Overview of the UFID
Intuition: Backdoor Attacks under a Causal Lens
Figure 1: Causal graph of clean and backdoored generation.
@@
To develop a backdoor detection algorithm for generative models, we first need to address a fundamental question: What distinguishes clean generation from backdoored generation, and how this distinction can be utilized in designing an effective detection algorithm? Motivated by the great potential of causal inference in deep learning, we propose to leverage causal inference as a new perspective to understand the distinct mechanisms underlying clean and backdoored generation processes. Specifically, we construct causal graphs to illustrate the comparison between the two processes, as shown in Figure 1.
A causal graph is a directed acyclic graph that illustrates the causal relationships among variables, where each node represents a variable and each edge represents a causal relationship. For simplicity, this figure only illustrates the unconditional diffusion model. However, the causal graph can be easily extended to the conditional diffusion model by substituting the input noise ($X_{T}$) with the input text ($T$).
@@
Clean Generation. As depicted in Figure 1(a), the generated image ($x_{0}^{c}$) is dependent on the input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$. This relationship is termed the causal path, denoted as $X_{T}^{c}\rightarrow X_{0}^{c}$. Consequently, adding a small Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ with a small weight $\alpha$ to $x_{T}^{c}$ results in a new input ${x^{\prime}}_{T}^{c}=x_{T}^{c}+\alpha\cdot\epsilon=\mathcal{N}({x^{\prime}}_{T}^{c};0,(1+\alpha)I)$, leading to a different generated image ${x^{\prime}}_{0}^{c}$, as shown in Figure 1(c).
@@
Backdoored Generation. As shown in Figure 1(b), a backdoor attack $A$ modifies an image $x_{T}$ by injecting a trigger $\delta$ and changing the image generation process towards the target image $x_{0}^{b}$, denoted as $X_{T}^{b}\leftarrow A\rightarrow X_{0}^{b}$, where $x_{T}^{b}=\delta+x_{T}^{c}$. This introduces a spurious path from $X_{T}^{b}$ to $X_{0}^{b}$, which lies outside the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$, thereby establishing and strengthening the erroneous correlation between the modified input noise and the target image. Consequently, generations of poisoned noise images are primarily influenced by this spurious path (Du et al. 2021; Zhang et al. 2023; Li et al. 2021), while the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$ plays a minor role, represented by a gray dotted line in Figure 1 (b). When an additional Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to the backdoored noise image $x_{T}^{b}$ with a small weight $\alpha$, the new backdoored input becomes ${x^{\prime}}_{T}^{b}\sim\mathcal{N}(\delta,(1+\alpha)I)$, which is a combination of a new noise image $x_{T}^{c}+\alpha\cdot\epsilon$ and the trigger pattern $\delta$. It can be interpreted as poisoning a new image $x_{T}^{c}+\alpha\cdot\epsilon$ with the trigger $\delta$. Hence, the generated image is still affected by the attack and lies within the domain of target images. In addition, the magnitude of the perturbation is controlled by a small weight $\alpha$ (e.g., 0.01), ensuring that the trigger pattern remains in the new input without being disrupted.
@@
In summary, for clean generation, a small perturbation significantly alters the output. However, triggers in backdoor samples tend to be robust features learned by neural network models. Consequently, minor perturbations of backdoor samples do not lead to substantial changes in the diffusion model’s generation results. The following theorems also validate our insights from causal analysis.
Lemma 1.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in Assumption 11 in the Appendix. Let input noise $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let $\hat{x_{0}}$ be the generated image for $x^{\prime}_{T}$ and the generated distribution for clean input $x_{T}^{c}$ be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I).$ (1)
The lemma implies that if we modify the variance of the input noise by multiplying it by $\rho^{2}$ (adding small Gaussian noise), the variance of the generated image is reduced to $\frac{1}{\rho^{2}}$ of the original.
Theorem 2.
Suppose the output domain of the diffusion model $f_{\theta}$ is Gaussian. Let $\mathcal{N}(x_{c},\sigma_{c})$ and $\mathcal{N}(x_{b},\sigma_{b})$ denote the distribution of clean generations and backdoor generations respectively. Let $N$ denote the image size. We assume that $\sigma_{c}\geq\sigma_{b}+\rho^{2}$. Given clean input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$, backdoor input noise $x_{T}^{b}=x_{T}^{c}+\delta\sim\mathcal{N}(\delta,I)$, and Lemma 1 if we perturb the clean input noises $x_{T}$ and backdoor input noise $x_{T}^{b}$ with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then for the resulted clean generations $\mathcal{N}(x^{\prime}_{c},\sigma^{\prime}_{c})$ and the backdoor generations $\mathcal{N}(x^{\prime}_{b},\sigma^{\prime}_{b})$, we have that $\sigma^{\prime}_{c}-\sigma^{\prime}_{b}\geq 1$.
Corollary 3.
@@
$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)\geq\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}>0.$ (2)
Theorem 2 and Corollary 3 imply that, after adding noise, the expected difference between the distance of generated clean images $\|x_{1}-x_{2}\|_{2}$ and the distance of the generated backdoor images $\|x_{3}-x_{4}\|_{2}$ is significantly larger than 0. This further sheds light on the intuition that the diversity of clean generations is significantly greater than backdoor generations after adding perturbations. Their proofs are provided in the Appendix.
Scenario 1: Unconditional Diffusion Models
@@
$y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I}.$ (3)
@@
Let $y_{i}=\{y_{i}^{j}|1\leq j\leq|\mathbb{M}|+1\}$ denote the generated batch. We can then determine whether the input query $x_{i}$ is a backdoored sample by inspecting the diversity of $y_{i}$.
Scenario 2: Conditional Diffusion Models